find 2 numbers such that the sum of the first and 3 times the second is 5 and the sum of second and two times the first is 8

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

find 2 numbers such that the sum of the first and 3 times the second is 5
a + 3b = 5
and the sum of second and two times the first is 8
2a + b = 8
:
multiply the 1st equation by 2, subtract from the 2nd equation
2a + 6b = 10
2a + b = 8
---------------Subtraction eliminates a, find b
0 + 5b = 2
b = 2/5 or .4 is the 2nd number
:
Find a using the 1st original equation
a + 3(.4) = 5
a + 1.2 = 5
a = 5 - 1.2
a = 3.8 is the 1st number
:
:
Check in the 2nd equation
2(3.8) + .4 = 8
7.6 + .5 = 8

Answer by Alan3354(69443)

find 2 numbers such that the sum of the first and 3 times the second is 5 and the sum of second and two times the first is 8
=======================
F + 3S = 5
2F + S = 8